OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0

OpenStudy (anonymous):

dumcow:can you go back to the other question when your finished with this one

OpenStudy (dumbcow):

x = -1.146, -7.854

OpenStudy (anonymous):

that isnt an option either

OpenStudy (anonymous):

\[\[(-9+3\sqrt{5})/2 , (-9-3\sqrt{5})/2\]\]